OpenStudy (anonymous):

a square garden ;lot mesasures 80sq fts. a second sq garden plot measures 320 sq fts. how many more feet of fence will the second garden plt require over the required for the firs garden plot? i got 80 ft is that right?

OpenStudy (anonymous):

Actually you are not it is \[48\sqrt{5}\] ft

OpenStudy (kropot72):

The length of each side of the 80 sq ft plot is as follows: \[\sqrt{80}\] The total length of fencing for the 80 sq ft plot is as follows: \[4\times \sqrt{80}=16\times \sqrt{5}\] The length of each side of the 320 sq ft plot is as follows: \[\sqrt{320}\] The total length of fencing for the 320 sq ft plot is as follows: \[4\times \sqrt{320}=32\times \sqrt{5}\] Extra length of fence needed for 320 sq ft plot is as follows: \[(32\times \sqrt{5})-(16\times \sqrt{5})=(16\times \sqrt{5})ft\]